OpenStudy (anonymous):

Anyone know how to do this problem? Given f(x) =x^3 and g(x) = 4x+3, Find g(f(3)).

jimthompson5910 (jim_thompson5910):

f(x)=x^3 f(3)=3^3 f(3)=27 --------------- g(f(3))=g(27) g(f(3))=4(27)+3 g(f(3))=108+3 g(f(3))=111

OpenStudy (anonymous):

Thank you for showing me the steps! I'll try to work it myself now. Thank you!

OpenStudy (anonymous):

In the first step, where do you get the number 3?:

jimthompson5910 (jim_thompson5910):

On the very inside, they want f(3). So I evaluated f(3) and got the number 27.

OpenStudy (anonymous):

ok. thank you